# A structure-aware framework for learning device placements on computation graphs

### v1

@@ Abstract
Existing approaches for device placement ignore the topological features of computation graphs and rely mostly on heuristic methods for graph partitioning. At the same time, they either follow a grouper-placer or an encoder-placer architecture, which requires understanding the interaction structure between code operations. To bridge the gap between encoder-placer and grouper-placer techniques, we propose a novel framework for the task of device placement, relying on smaller computation graphs extracted from the OpenVINO toolkit using reinforcement learning. The framework consists of five steps, including graph coarsening, node representation learning and policy optimization. It facilitates end-to-end training and takes into consideration the directed and acyclic nature of the computation graphs. We also propose a model variant, inspired by graph parsing networks and complex network analysis, enabling graph representation learning and personalized graph partitioning jointly, using an unspecified number of groups. To train the entire framework, we utilize reinforcement learning techniques by employing the execution time of the suggested device placements to formulate the reward. We demonstrate the flexibility and effectiveness of our approach through multiple experiments with three benchmark models, namely Inception-V3, ResNet, and BERT. The robustness of the proposed framework is also highlighted through an ablation study. The suggested placements improve the inference speed for the benchmark models by up to $58.2\%$ over CPU execution and by up to $60.24\%$ compared to other commonly used baselines.
@@ 1 Introduction
The ability of intelligent agent systems to perceive and accurately interpret complex environments is crucial for artificial intelligence (AI). Recently, the AI community explores a variety of transformer architectures and foundation models (FMs) [28, 32]. FMs have allowed both academia and industry to perform several data-demanding tasks, ranging from image and text analysis to multi-modal content generation and human-like visual perception [20, 22, 11, 4]. This is achievable due to the self-supervised nature of the FMs, their ability to easily generalize and the large amounts of data available online [31]. The aforementioned properties make FMs distinguishable in specific general pre-training tasks such as next-word prediction, compared to traditional ML architectures that use supervised learning [2].
This recent surge in using large deep learning models (both training/inference) has led to increasing the demand for computing power which is projected to grow even more in the next few years [18]. This is due to the need for more advanced model training processes and the continuous expansion of model parameters [6, 28]. It is clear that as FMs become increasingly complex, they demand vast computational resources not only for training but also for fine-tuning and inference tasks. This surge in computational demand underscores the necessity for managing the available hardware more effectively.
@@
In the beginning, device placement has been the main responsibility of human experts [23]. Engineers with a substantial level of expertise and diligence were responsible for allocating each part of a model to the best-suited device. However, this rigorous task can be daunting, considering the rapid advancement in hardware, which leads to a serious increase in development time, bug fixing, and code optimization [3, 21, 25]. Deep reinforcement learning (DRL) has recently been proposed to provide effective device placements with full automation [23, 1, 33]. Two different DRL architectures for device placement currently exist in the literature: the ‘grouper-placer’ model that reduces the action space by merging operations into groups; and the ‘encoder-placer’ that encodes the features of the operations to capture the topological properties of the computational graph [17].
@@
Although existing approaches have been successful, there is still space for improvement. In particular, they demonstrate several shortcomings. To begin with, they disregard the directed and acyclic nature of computation graphs. Furthermore, they either follow a grouper-placer or an encoder-placer architecture. In addition, most of them are not designed to train all of their components simultaneously in an end-to-end fashion. Additionally, they fail to capture higher-order interactions among the operations of a computation graph. Most notably, they make use of large, fine-grained computation graphs, thereby exhibiting slow convergence and demanding a higher number of iterations during the learning process.
@@
Considering the limitations mentioned above, this paper proposes a framework to optimize the task of device placement based on smaller, coarsened computation graphs produced by the OpenVINO toolkit. Our framework consists of five steps, where a neural network model is first converted into a computation graph. Then, local and global structural features as well as positional, node-specific and features from fractal analysis are extracted to compose the initial node feature vectors. Then, graph representation learning, graph partitioning and pooling are learned jointly, facilitating the fusion of the grouper-placer and encoder-placer models. Finally, we use the execution time of the suggested device placements as a reward to train the entire framework. In contrast with existing approaches, our framework allows for encoding and grouping operations of a computation graph jointly in an end-to-end fashion. Along with the proposed framework, through one of the variant models, we introduce a novel method tailored to computation graphs, for jointly learning node embeddings and performing personalized graph partitioning with an unspecified number of groups for further coarsening. The effectiveness and robustness of the proposed approach are demonstrated through multiple experiments with different benchmark models and a detailed ablation study. The list of benchmark models includes Inception-V3, ResNet and BERT. The proposed framework demonstrates a performance improvement of up to $58.2\%$ over CPU execution and up to $60.24\%$ compared to other commonly used baselines.
@@
  • Our framework first illustrates the combination of features from fractal analysis, positional encodings, and node-specific features for the task of device placement through a model variant, and discusses the impact of incorporating different properties on the model.
@@ 2 Proposed framework
In this section, we introduce our framework, HSDAG. It consists of five steps as shown in Figure 1. Briefly, we first convert a neural network model into a computation graph. Then, we extract features for each node and edge of the computation graph. The next step enriches these features via graph representation learning techniques, while it learns how to partition and pool computation graphs simultaneously. The learned features and groups of nodes are utilized to train a stochastic policy on the device placement problem. Finally, it can be employed to assign each node of a computation graph to the most appropriate device, resulting in minimum inference time. The entire process is end-to-end trainable.
@@ 2.1 Problem Formulation
Problem setup. We focus on the problem of device placement in a heterogeneous computing system. Our main goal is to assign each part of a computation graph to the most suitable hardware device, in such a way that the overall execution time during the inference of the model is minimized. Formally, given a computation graph $G$, we learn a policy $\pi:G\rightarrow P$ that assigns a placement $p$ for all $v\in G$ such that
@@ 2.3 Feature extraction
Global structural features. Relying solely on local features might miss capturing important global properties of the network. To capture the multi-scale structural properties of the network, we calculate the fractal dimension for each node $v\in V$. The fractal dimension $D(v)$ of a node $v$ is computed based on the mass distribution. Given the set of distances $\{r_{1},r_{2},\ldots,r_{m}\}$ from node $v$ to other nodes in the network, the fractal dimension $D(v)$ is calculated as follows:
@@ 3.1 Benchmarks
To evaluate our approach we use the computation graphs created from three popular benchmarks: (1) Inception-V3: The Inception-V3 architecture [27] is extensively employed for image recognition and visual feature extraction [12]. This neural network consists of multiple blocks, each comprising various branches of convolutional and pooling layers. These branches are capable of parallel execution and are concatenated to form inputs for the subsequent block. However, the depth of the network limits this parallelism since later blocks must wait for the completion of earlier ones; (2) ResNet: ResNet [10] is a widely-used model for image classification. It consists of multiple convolutional layers and uses residual connections to reduce the effects of the vanishing gradient problem. We use the ResNet-50, which is a 50-layer convolutional neural network; (3) BERT: BERT [6] is a language model relying on the transformer architecture. It pre-trains deep bidirectional representations on unlabeled data jointly. It can be used as the base to fine-tune models for a list of tasks, including question answering and language inference. Several versions of the BERT model exist. In this paper, we use the base uncased version. Important statistics and a more detailed description of the benchmark models are available in Table 1 and Appendix E.
@@ 3.3 Baseline comparison
OpenVINO-CPU. This baseline method lets the OpenVINO optimization toolkit decide whether the entire computation graph should be assigned to CPU or GPU but we set CPU as the first preference.
OpenVINO-GPU. A baseline similar to OpenVINO-CPU but we set GPU as the first preference.
Placeto [1]. It employs GNNs to learn features for any computation graph as proposed in [1]. Therefore it enables the transfer of a learned device placement policy to new computation graphs without further re-training.
@@ 3.4 Ablation studies
To understand the impact of the components, steps and configurations of HSDAG, we conduct additional experiments. Results are shown in Table 3. Various modifications to the framework were tested, such as removing graph structural features, output shape features, and node IDs. Overall, the results indicate that each of these components plays a significant role in achieving optimal performance.
No graph structural features (w/o graph structural features). For each node of $v$ the computation graph $G$, we ignore features from fractal analysis, in-degree, and out-degree. Removing graph structural features clearly impacts the framework’s ability to capture the global and local structural information of the computation graph. The results from this ablation show a decrease in performance, although not as important as other feature removals. For example, the inference time speedup on Inception-V3 drops to $14.8\%$ from $17.9\%$ when these features are excluded. This indicates the importance of these features in capturing the hierarchical and interconnected nature of computation graphs.
No output shape features (w/o output shape). We assume that the output data shape of each operation reflects the computation requirement for the corresponding operation. We do not include the output data shape in the node feature vector to test its effectiveness. We observe a significant decrease in performance. The speedup for Inception-V3 drops from $17.9\%$ to $8.59\%$. Similar reductions are also observed on ResNet and BERT. This suggests that the features related to the output shape are crucial for understanding the computational load associated with each node in the graph. Without this information, the framework’s ability to optimize device placement is compromised, leading to less efficient computation graph execution.
No node ID (w/o node ID). In this case, we do not use the node ID to encode the topological sequence of the node $v$ in a given computation graph $G$. Omitting information about the node ID results in a significant performance drop. The speedup for Inception-V3 is reduced to $8.59\%$. This highlights the critical role of node IDs in preserving the order and dependencies within the computation graph. The framework lacks essential information about the execution order of operations, leading to suboptimal device placement and reduced overall efficiency.
@@ 4 Conclusion
In this work, we introduce a flexible framework for the task of device placement. Our framework relies on smaller computation graphs is divided into five steps such as graph coarsening, node representation learning and policy optimization using reinforcement learning. It supports end-to-end model training and is aware of the directed and acyclic nature of the computation graphs under consideration. Additionally, a model variant has been proposed that facilitates graph representation learning and personalized graph partitioning jointly with an unspecified number of node clusters. The experimental results highlighted that our framework is flexible, and robust and mitigates the shortcomings of the grouper-placer and encoder-placer models by capitalizing on the best aspects of the two worlds. The suggested placements improve the inference speed for the benchmark models compared to widely used baselines.
Limitations. The measurement did not consider the temperature change of the environment of the system and surroundings. During our experiments, we allocate the CPU for both experiment and policy due to the limitation of the setup. At the same time, iGPU is not considered throughout the experiment because we consider it to be always slower than both CPU and dGPU. We attempted to obtain the he source codes for the baseline methods, but they were not made available from the corresponding authors. As such, we implemented and reproduced the baseline methods with our best effort from their published papers.
Appendix A Code and data availability
@@ C.1 Message-Passing neural networks
Appendix D Graph neural networks for graph information encoding
Many different models can be employed as the first GNN component to learn jointly embeddings and graph partitioning algorithms as described in Section 2.4. We present some of the most popular choices below. Although the selected model is not required to be capable of capturing information about a graph, we empirically observed that the utilization of a GNN is beneficial for the overall framework.
D.1 Graph convolutional network
The graph convolutional operator, as proposed in [16], for each convolutional layer is defined as follows:
$\mathbf{Z}=\sigma(\hat{\mathbf{D}}^{-1/2}\hat{\mathbf{A}}\hat{\mathbf{D}}^{-1/2}\mathbf{X}\mathbf{W})\in\mathbb{R}^{|V|\times F}$ (19)
where $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}\in\{0,1\}^{|V|\times|V|}$ denotes the adjacency matrix with self-loops and $\hat{\mathbf{D}}_{ii}=\sum_{j=0}\hat{\mathbf{A}}_{ij}$ is the corresponding diagonal degree matrix, $\mathbf{W}\in\mathbb{R}^{C\times F}$ is a matrix of learnable parameters, $\mathbf{X}\in\mathbb{R}^{|V|\times C}$ is a matrix with the input features of each node, $\sigma(\cdot)$ denotes an activation function such as $\text{ReLU}(\cdot)=\max(0,\cdot)$ or $\text{PReLU}(\cdot)=\max(0,\cdot)+a\cdot\min(0,\cdot)$, $F$ and $C$ the size of the output and input embedding, respectivelly.
D.2 Graph attention network
The Graph Attention (GAT) [29] layer builds upon the self-attention mechanism, where each node $u$ receives a message from its neighbor $\mathcal{N}(u)$ using a shared attentional mechanism $a:\mathbb{R}^{F^{\prime}}\times\mathbb{R}^{F^{\prime}}\rightarrow\mathbb{R}$, where $F^{\prime}$ is the size of the output feature vector. The attention coefficients during the self-attention process are calculated as follows:
$e_{u,v}=\alpha(\mathbf{W}h_{u},\mathbf{W}h_{v})$ (20)
where $h_{u}\in\mathbb{R}^{F}$ and $h_{v}\in\mathbb{R}^{F}$ are the input features of node $u$ and $v$, respectively. The attention coefficients $e_{u,v}$ indicate the importance of node $v$’s features to node $u$. It then constructs a final representation matrix $\mathbf{Z}\in\mathbb{R}^{|V|\times F^{\prime}}$ using an aggregation operator. Thus, a GAT network is able to understand hidden relationships among the nodes of a computation graph $G$. For each node $u\in V$ of a graph $G$ each GAT layer is formulated as follows:
$\mathbf{z}_{u}=a_{u,u}\mathbf{W}\mathbf{h}_{u}+\sum_{v\in\mathcal{N}(u)}a_{u,v}\mathbf{W}\mathbf{h}_{v}\in\mathbb{R}^{|V|\times F^{\prime}}$ (21)
where $\mathbf{W}\in\mathbb{R}^{F^{\prime}\times F}$ is a trainable weight matrix. These attention parameters are calculated via a dense neural network architecture which employs a $\text{softmax}(e_{u,v})=\frac{\exp({e_{u,v}})}{\sum_{k\in\mathcal{N}(u)\exp(e_{u,k})}}$ activation function:
$a_{u,v}=\text{softmax}(e_{u,v})=\frac{\exp({e_{u,v}})}{\sum_{k\in\mathcal{\hat{N}}(u)\exp(e_{u,k})}}=\frac{\exp(\text{LeakyReLU}(\mathbf{a}^{T}[\mathbf{W}\mathbf{h}_{u}\oplus\mathbf{Wh}_{v}]))}{\sum_{k\in\mathcal{\hat{N}}(u)}\exp(\text{LeakyReLU}(\mathbf{a}^{T}[\mathbf{W}\mathbf{h}_{u}\oplus\mathbf{Wh}_{k}]))}$ (22)
where $\mathbf{a}$ is a trainable attention vector, $\oplus$ denotes the concatenation operation and $\mathcal{\hat{N}}(u)=\mathcal{N}(u)\cup\{u\}$.
D.3 GraphSAGE
In GraphSAGE [8], given a graph $G=(V,E)$, where $V$ is the set of nodes and $E$ is the set of edges, the goal is to learn a function $f:V\rightarrow\mathbb{R}^{F}$ that maps each node to an $F$-dimensional embedding space.
D.3.1 Neighborhood sampling
For each node $v\in V$, GraphSAGE samples a fixed-size set of neighbors $\mathcal{N}_{v}\subseteq\{u\in V:(u,v)\in E\}$. This can also be expressed as $N_{v}=\text{SampleNeighbors}(v,S)$ where $S$ denotes the sample size.
D.3.2 Aggregation function
GraphSAGE defines an aggregation function AGGREGATEk at layer $k$, which aggregates features from the neighbors of node $v$. Let $\mathbf{h}_{u}^{(k)}$ represent the embedding of node $u$ at layer $k$. The aggregation process is given by:
$\mathbf{h}_{\mathcal{N}_{v}}^{(k)}=\text{AGGREGATE}_{k}(\{\mathbf{h}_{u}^{(k-1)}|u\in\mathcal{N}_{v}\})$ (23)
Common aggregation functions include the mean, LSTM and pooling aggregator:
$\mathbf{h}_{\mathcal{N}_{v}}^{(k)}=\frac{1}{|\mathcal{N}_{v}|}\sum_{u\in\mathcal{N}_{v}}\mathbf{h}_{u}^{(k-1)},\mathbf{h}_{\mathcal{N}_{v}}^{(k)}=\text{LSTM}_{k}(\{\mathbf{h}_{u}^{(k-1)}|u\in\mathcal{N}_{v}\}),\mathbf{h}_{\mathcal{N}_{v}}^{(k)}=\max_{u\in\mathcal{N}_{v}}(\sigma(\mathbf{W}\mathbf{h}_{u}^{(k-1)}+\mathbf{b}))$ (24)
D.3.3 Update function
Following aggregation, the node’s representation is updated by concatenating its previous representation with the aggregated neighborhood vector, hollowed by a linear transformation and nonlinearity:
$\mathbf{h}_{v}^{(k)}=\sigma(\mathbf{W}^{(k)}\cdot[\mathbf{h}_{v}^{(k-1)}||\mathbf{h}_{\mathcal{N}_{v}}^{(k)}])$ (25)
where $\sigma$ is a nonlinear activation function such as ReLU, $\mathbf{W}^{(k)}$ is a trainable weight matrix, and [||][$\cdot$|$\cdot$][ ⋅ | | ⋅ ] denotes vector concatenation.
D.4 Varational Graph Autoencoder
A Variational Graph Autoencoder (VGAE) consists of two models, an encoder and a decoder model [15]. The main goal of a VGAE model is to train a probabilistic decoder model $p_{\theta}$, from which we can sample realistic graphs $\hat{A}\sim p_{\theta}(\mathbf{A|Z})$ by conditioning on a latent variable $\mathbf{Z}\sim p_{\theta}(Z)$. More specifically:
$q_{\theta}(\mathbf{Z}|G)=\prod_{i=1}^{N}q_{\theta}(\mathbf{z}_{i}|G)$ (26)
where $G=(\mathbf{V},\mathbf{X},\mathbf{A})$, $q_{\theta}(\mathbf{z}_{i}|G)=\mathcal{N}(\mathbf{z}_{i}|\mathbf{\mu}_{i},\text{diag}(\sigma^{2}_{i}))$, $\mathbf{\mu}=\text{GNN}_{\mathbf{\mu}}(G)\in\mathbb{R}^{|V|\times F}$ is the matrix of mean vectors $\mathbf{\mu}_{i}$, $\log\mathbf{\sigma}=\text{GNN}_{\mathbf{\sigma}}(G)\in\mathbb{R}^{|V|\times F}$ and $F$ is the size of the embedding space. To calculate $\mathbf{\mu}$ and $\log\mathbf{\sigma}$ any graph neural network can be employed, including GraphSAGE, GCN and Graph transformers.
Given a set of training graphs $\mathcal{G}=\{G_{1},G_{2},...,G_{n}\}$ a VGAE model can be trained using the variational lower bound $\mathcal{L}$ by minimizing the evidence likelihood lower bound (ELBO) considering as follows:
$\mathcal{L}=\sum_{G_{i}\in\mathcal{G}}\mathbb{E}_{q_{\theta}(\mathbf{Z}|G_{i})}[\log p_{\theta}(G_{i}|\mathbf{Z})]-\text{KL}[q_{\theta}(\mathbf{Z}|G_{i})||p_{\theta}(\mathbf{Z})]$ (27)
where KL is the Kullback-Leibler divergence between $q_{\theta}(\cdot)$ and $p_{\theta}(\cdot)$ using a Gaussian prior $p_{\theta}(\mathbf{Z})=\prod_{i}p_{\theta}(\mathbf{z_{i}})=\prod_{i}\mathcal{N}(\mathbf{z_{i}}|0,\mathbf{I})$.
D.5 Graph Autoencoder
A non-probabilistic version of VGAE is the Graph Autoencoder model (GAE) [15]. In GAE, the embeddings $\mathbf{Z}\in\mathbb{R}^{|V|\times F}$ are calculated by taking into consideration the reconstruction of the adjacency matrix $\hat{A}\in\mathbb{R}^{|V|\times|V|}$ exists such that $\hat{A}=\sigma(\mathbf{ZZ}^{T})$ and the embeddings $\mathbf{Z}$ can be generated from any neural network model, preferably a graph neural network $\mathbf{Z}=\text{GNN}(G)$.
Both the VGAE and GAE models can be a great choice as they allow for pre-training over a predefined general task such as link prediction.
Appendix E Benchmark models
E.1 Inception-V3

### v2

@@ Abstract
Computation graphs are Directed Acyclic Graphs (DAGs) where the nodes correspond to mathematical operations and are used widely as abstractions in optimizations of neural networks. The device placement problem aims to identify optimal allocations of those nodes to a set of (potentially heterogeneous) devices. Existing approaches rely on two types of architectures known as grouper-placer and encoder-placer, respectively. In this work, we bridge the gap between encoder-placer and grouper-placer techniques and propose a novel framework for the task of device placement, relying on smaller computation graphs extracted from the OpenVINO toolkit. The framework consists of five steps, including graph coarsening, node representation learning and policy optimization. It facilitates end-to-end training and takes into account the DAG nature of the computation graphs. We also propose a model variant, inspired by graph parsing networks and complex network analysis, enabling graph representation learning and jointed, personalized graph partitioning, using an unspecified number of groups. To train the entire framework, we use reinforcement learning using the execution time of the placement as a reward. We demonstrate the flexibility and effectiveness of our approach through multiple experiments with three benchmark models, namely Inception-V3, ResNet, and BERT. The robustness of the proposed framework is also highlighted through an ablation study. The suggested placements improve the inference speed for the benchmark models by up to $58.2\%$ over CPU execution and by up to $60.24\%$ compared to other commonly used baselines.
@@ 1 Introduction
The ability of intelligent agent systems (IASs) and cyber-physical systems (CPSs) to perceive and accurately interpret complex environments is crucial for artificial intelligence (AI). Recently, there has been a remarkable progress in machine learning and AI, due to the wide adoption of the transformer architecture and foundation models (FMs) [26, 30]. FMs have allowed both academia and industry to perform several data-demanding tasks, ranging from image and text analysis to multi-modal content generation and human-like visual perception [18, 20, 11, 4]. This is achievable due to the self-supervised nature of the FMs, their ability to easily generalize and the large amounts of data available online [29]. The aforementioned properties make FMs distinguishable in specific general pre-training tasks such as next-word prediction, compared to traditional ML architectures that use supervised learning [2]. This recent surge in using large FMs has led to increased demand for computing power, which is projected to grow even more in the next few years [16]. This is due to the need for more advanced model training processes and the continuous expansion of model parameters [26, 6]. It is clear that as FMs become increasingly complex, they demand vast computational resources not only for training but also for fine-tuning and inference tasks. This surge in computational demand underscores the necessity for managing the available hardware more effectively.
@@
Early on device placement has been the main responsibility of human experts [21]. Engineers with a substantial level of expertise and diligence were responsible for allocating each part of a model to the best-suited device. However, this rigorous task can be daunting, considering the rapid advancement in hardware, which leads to a serious increase in development time, bug fixing, and code optimization [3, 19, 23]. Deep reinforcement learning (DRL) has recently been proposed to provide effective device placements with full automation [31, 1, 21]. Two different DRL architectures for device placement currently exist in the literature: the ‘grouper-placer’ model that reduces the action space by merging operations into groups; and the ‘encoder-placer’ that encodes the features of the operations to capture the topological properties of the computational graph [15, 31].
@@
Although existing approaches have been successful, there is still space for improvement. In particular, they demonstrate several shortcomings. To begin with, they disregard the directed and acyclic nature of computation graphs [22]. Furthermore, they either follow a grouper-placer or an encoder-placer architecture [22, 21]. In addition, most of them are not designed to train all of their components simultaneously in an end-to-end fashion and they fail to capture higher-order interactions among the operations of a computation graph. Finally, they make use of large, fine-grained computation graphs, thereby exhibiting slow convergence and demanding a higher number of iterations during the learning process [15].
@@
Considering the limitations mentioned above, this paper proposes a framework to optimize for device placement based on smaller, coarsened computation graphs produced by the OpenVINO toolkit. Our framework consists of five steps: First, the neural network model is converted into a computation graph. Then, local and global structural features as well as positional, node-specific and fractal features are extracted to compose the initial node feature vectors. Following that, graph representation learning, graph partitioning and pooling are learned jointly, facilitating the fusion of the grouper-placer and encoder-placer models. Finally, we use the execution time of the suggested device placements as a reward to train the entire framework. In contrast with existing approaches, our framework allows for encoding and grouping operations of a computation graph jointly in an end-to-end fashion. Along with the proposed framework, through one of the variant models, we introduce a novel method tailored to computation graphs, for jointly learning node embeddings and performing personalized graph partitioning with an unspecified number of groups for further coarsening. The effectiveness and robustness of the proposed approach are demonstrated through multiple experiments with different benchmark models and a detailed ablation study.
@@
  • Our framework is the first of its kind that encodes features from multifractal analysis, positional encodings, and node-specific features for the task of device placement through a model variant, and discusses the impact of incorporating different properties on the model.
@@ 2 Proposed framework
In this section we introduce our framework titled Hierarchical Structure-Aware Device Assignment Graph (HSDAG). It consists of five steps as shown in Figure 1. Briefly, we first convert a neural network model into a computation graph. Then, we extract features for each node and edge of the computation graph. The next step enriches these features via graph representation learning techniques and simultaneously learns how to partition and pool the graphs. The learned features and groups of nodes are utilized to train a stochastic policy, which we use for assigning each node of a graph to the most appropriate device. We train the entire pipeline end-to-end with the objective of minimizing the inference time of the proposed device placement.
@@ 2.1 Problem Formulation
Problem setup. We focus on the problem of device placement in a heterogeneous computing system. Our goal is to assign each part of a computation graph to the most suitable device, such that the overall execution time during the inference of the model is minimized. Formally, given a computation graph $G$, we learn a policy $\pi:G\rightarrow P$ that assigns a placement $p$ for all $v\in G$ such that
@@ 2.3 Feature extraction
Global structural features. Relying solely on local features might miss capturing important global properties of the network. To capture the multi-scale structural properties of the network, we calculate the fractal dimension for each node $v\in V$. The fractal dimension $D(v)$ [28, 30] of a node $v$ is computed based on the mass distribution. Given the set of distances $\{r_{1},r_{2},\ldots,r_{m}\}$ from node $v$ to other nodes in the network, the fractal dimension $D(v)$ is calculated as follows:
@@ 3.1 Benchmarks
To evaluate our approach we use the computation graphs created from three popular benchmarks: (1) Inception-V3: The Inception-V3 architecture [25] is extensively employed for image recognition and visual feature extraction [12]. This neural network consists of multiple blocks, each comprising various branches of convolutional and pooling layers. These branches are capable of parallel execution and are concatenated to form inputs for the subsequent block. However, the depth of the network limits this parallelism since later blocks must wait for the completion of earlier ones; (2) ResNet: ResNet [10] is a widely-used model for image classification. It consists of multiple convolutional layers and uses residual connections to reduce the effects of the vanishing gradient problem. We use the ResNet-50, which is a 50-layer convolutional neural network; (3) BERT: BERT [6] is a language model relying on the transformer architecture. It pre-trains deep bidirectional representations on unlabeled data jointly. It can be used as the base to fine-tune models for a list of tasks, including question answering and language inference. Several versions of the BERT model exist. In this paper, we use the base uncased version. Important statistics and a more detailed description of the benchmark models are available in Table 1 and Appendix D.
@@ 3.3 Baseline comparison
  3. 3.
    OpenVINO-CPU. This baseline method lets the OpenVINO optimization toolkit decide whether the entire computation graph should be assigned to CPU or GPU, with CPU set as the first preference.
  4. 4.
    OpenVINO-GPU. A baseline similar to OpenVINO-CPU with GPU set as the first preference.
  5. 5.
    Placeto [1]. It uses GNNs to learn features for any computation graph as proposed in [1]. Therefore it enables the transfer of a learned device placement policy to new computation graphs without further re-training.
  6. 6.
@@ 3.4 Ablation studies
To understand the impact of the components, steps and configurations of HSDAG, we conduct an ablation study. Various modifications to the framework were tested, such as removing graph structural features, output shape features, and node IDs. Overall, the results are shown in Table 3. indicate that each of these components plays a significant role in achieving optimal performance.
No graph structural features. For each node of $v$ in the computation graph $G$, we ignore features from fractal analysis, in-degree, and out-degree. Removing graph structural features clearly impacts the framework’s ability to capture the global and local structural information of the computation graph. The results from this ablation show a decrease in performance, although not as important as other feature removals. For example, the inference time speedup on Inception-V3 drops to $14.8\%$ from $17.9\%$ when these features are excluded. This indicates the importance of these features in capturing the hierarchical and interconnected nature of computation graphs.
No output shape features. We assume that the output data shape of each operation reflects the computation requirement for the corresponding operation. We do not include the output data shape in the node feature vector to test its effectiveness. We observe a significant decrease in performance. The speedup for Inception-V3 drops from $17.9\%$ to $8.59\%$. This suggests that the features related to the output shape are crucial for understanding the computational load associated with each node in the graph. Without this information, the framework’s ability to optimize device placement is compromised, leading to less efficient computation graph execution.
No node ID. In this case, we do not use the node ID to encode the topological sequence of the node $v$ in a given computation graph $G$. Omitting information about the node ID results in a significant performance drop. The speedup for Inception-V3 is reduced to $8.59\%$. This highlights the critical role of node IDs in preserving the order and dependencies within the computation graph. The framework lacks essential information about the execution order of operations, leading to sub-optimal device placement and reduced overall efficiency.
@@
3.5 Downstream Model Performance and Runtime Complexity
As a sanity check for our method, we show the performance on downstream tasks is not affected. Theoretically, since the end-to-end training pipeline itself does not change, we do not expect any impact on the performance. Nonetheless, we show experimental results on 3 exemplar cases to empirically verify statement.
Inception-V3: We performed image classification inference on images depicting Samoyed dogs. All the parameters are directly derived from the torchvision pre-trained model. We did not change any configuration on the data type of the model. The classification accuracy of Inception-V3 using the best device placement is 82.77%. For the GPU-only experiments the classification accuracy is 82.72%. For the CPU-only experiments the classification accuracy is 82.33%.
ResNet: Similarly, we performed image classification inference using the ResNet model on the same dataset. The classification accuracy with the best device placement is 45.37%. For the GPU-only experiments the classification accuracy is 45.37%. For the CPU-only experiments the classification accuracy is 45.44%.
BERT: We evaluated the performance of the BERT model using the output embeddings from the different device placements. We calculated their mean squared error, cosine similarity and Euclidean distance (MSE: the lower the better, Cosine Similarity: the higher the better, euclidean distance: the lower the better). The results are shown on Table 5.
Table 4: BERT performance on downstream tasks.
Comparison MSE CS L2 norm
CPU vs GPU 3.049e-05 0.999 0.432
CPU vs HSDAG 6.819e-07 0.999 0.064
GPU vs HSDAG 3.174e-05 0.999 0.441
Table 5: Empirical runtime complexity comparison.
Model Inception-V3 ResNet BERT
Placeto 2808s 1162s 4512s
RNN-based 3706s 1212s OOM
HSDAG 2454s 1047s 2765s
These experiments empirically demonstrate that HSDAG does not affect the performance of the model in the downstream tasks. All models have similar performance regardless of the running device (e.g. CPU, GPU or heterogeneous device). Finally, we also conducted an empirical runtime complexity estimation in order to show HSDAG’s superiority in terms of execution time. We compare against the Placeto and the RNN-based device placement methods for the same 3 models. The results are shown in Table 5.
@@ 4 Conclusion
We introduce a flexible framework for the task of device placement. Our framework relies on smaller computation graphs and is divided into five steps such as graph coarsening, node representation learning and policy optimization using reinforcement learning. It supports end-to-end model training and is aware of the directed and acyclic nature of the computation graphs under consideration. Additionally, we propose a model variant that facilitates graph representation learning and personalized graph partitioning jointly with an unspecified number of node clusters. The experimental results highlighted that our framework is flexible, and robust and mitigates the shortcomings of the grouper-placer and encoder-placer models by capitalizing on the best aspects of the two worlds. The suggested placements improve the inference speed for the benchmark models compared to widely used baselines. One interesting future work direction is to explore different RL problem formulations as well as different reward structures, such as the incremental rewards used in [1]. Another direction is to study the interplay of the reward and generalizability as well as potentially using reward models rather than measuring reward. This could unlock a more efficient algorithm as the current setup relies on measuring the inference latency, which has practical limitations.
Limitations. The latency measurements did not consider the temperature change of the environment of the system and surroundings. During our experiments, we allocate the CPU for both experiment and policy due to the limitation of the setup. At the same time, iGPU is not considered throughout the experiment because we consider it to be always slower than both CPU and GPU. We attempted to obtain the source code for the baseline methods, but they were not made available from the corresponding authors. As such, we implemented and reproduced the baseline methods with our best effort from their published papers.
Acknowledgements
The University of Southern California acknowledges the support by the U.S. Army Research Office (ARO) under Grant No. W911NF-23-1-0111, the National Science Foundation (NSF) under the Career Award CPS-1453860, CCF-1837131, MCB-1936775, CNS-1932620 and the NSF award No. 2243104 under the Center for Complex Particle Systems (COMPASS), the Defense Advanced Research Projects Agency (DARPA) Young Faculty Award and DARPA Director Award under Grant Number N66001-17-1-4044, National Institutes of Health (NIH) under grant R01 AG 079957, an Intel faculty award and a Northrop Grumman grant. The views, opinions, and/or findings in this article are those of the authors and should not be interpreted as official views or policies of the Department of Defense, the National Institutes of Health, or the National Science Foundation.
Appendix A Code availability
The source code is available at https://github.com/hping666/HSDAG.
@@ C.1 Message-Passing neural networks
Appendix D Benchmark models
D.1 Inception-V3